# Unlocking the Potential of Model Calibration in Federated Learning

### v2

@@ 4.2 Non-Uniform Calibration for Federated Learning (NUCFL)
Motivation: Limitation of uniform penalty. In Equation (5), we promote calibration during federated local training by targeting the confidence-accuracy gap in individual local models via auxiliary calibration loss. The main aim of FL is to train a global model capable of effectively handling the global data distribution. Therefore, we aim for calibration to account for the heterogeneity inherent in FL setups, marking a departure from calibration approaches used in centralized learning. Direct application of auxiliary loss calibration methods without adapting to FL characteristics would result in uniform calibration weights, such that $\beta_{1}=\beta_{2}=...=\beta_{m}$. This leads to local models being calibrated solely to their respective datasets $D_{m}$, with uniformly small weights biasing calibration toward local heterogeneity, and uniformly large weights potentially neglecting accuracy improvements from classification. Such discrepancies can result in a global model that, while comprising well-calibrated local models, does not achieve optimal performance across the broader distribution. This fixed calibration penalty approach, therefore, risks neglecting the broader global calibration needs necessary for optimal performance across the entire global distribution.
@@
Non-uniform penalty design (NUCFL). Given that each client receives the global model from the previous round as their starting point for each new federated round, this offers an opportunity to indirectly adjust local calibration to mirror global model characteristics. To leverage this, we propose non-uniform calibration for federated learning (NUCFL), which provides a dynamic calibration penalty to local models to better reflect global calibration needs. Drawing inspiration from the concept of similarity characterization in FL techniques Sahu et al. [2018]; Tan et al. [2023], we hypothesize that a local model closely resembling the global model is likely to represent global characteristics well, suggesting that the penalty appropriately reflects the calibration needs of the global model. Conversely, a dissimilar/heterogeneous local model suggests a focus on local objectives (e.g., to improve accuracy) at the expense of global alignment. By adjusting penalties at each local epoch based on the similarity between local and global models, NUCFL ensures that local training enhances both the calibration and accuracy of the global model. As illustrated in Fig. 2, we specifically set
@@
where $\delta^{(t-1)}$ is the accumulated gradient from the previous round $t-1$, representing the received global model, while $\delta_{m}^{(t,e)}={w}^{(t-1)}-{w}^{(t,e)}_{m}$ calculates the gradient for client $m$ at local epoch $e$ during round $t$. The function sim($\cdot$) measures similarity between local and global models, generating values between 0 and 1. Specifically, it can use general cosine similarity or more advanced measures like Centered Kernel Alignment (CKA) Kornblith et al. [2019].
For instance, if the learned local model is similar to the received global model, indicated by a higher similarity score, it suggests that the local model’s characteristics align well with those of the global model. In such cases, the calibration penalty applied to the local model is increased to reflect the global calibration needs. Otherwise, i.e., if the similarity score is low, we choose a smaller penalty and instead focus on improving the accuracy. Based on this non-uniform penalty method tailored to FL, our scheme strategically improves model calibration as well as the accuracy.
Remark 1. NUCFL can integrate any auxiliary-based calibration loss and adapt to any FL algorithm regardless of aggregation method or local loss function. To the best of our knowledge, this is the first work to address model calibration specifically for FL, introducing a tailored calibration method. Similar to most calibration research Lin et al. [2017]; Müller et al. [2019]; Brier [1950]; Kumar et al. [2018]; Mukhoti et al. [2020]; Liu et al. [2021]; Liang et al. [2020]; Hebbalaguppe et al. [2022b; a], our work is an empirical study that explores the practical impacts of our methods. This area of research remains challenging, particularly in providing theoretical guarantees for calibration methods, even in centralized settings; we instead substantiate the effectiveness of our method through extensive experiments. We will demonstrate how our strategic approach effectively calibrates FL in Section 5.2.
@@ 5.1 Experimental Setup
Dataset and model. We conduct experiments using four image classification datasets commonly utilized in FL research Caldas et al. [2018]; McMahan et al. [2017]; Mohri et al. [2019]: MNIST LeCun et al. [1998], FEMNIST Cohen et al. [2017], CIFAR-10 Krizhevsky [2009], and CIFAR-100 Krizhevsky [2009]. For the MNIST and FEMNIST datasets, we use a CNN LeCun & Bengio [1998] with two 5x5 convolution layers, each followed by 2x2 max pooling, and fully connected layers with ReLU activation. We use AlexNet Krizhevsky et al. [2012] with five convolution layers for the CIFAR-10/100 datasets.
FL data distribution. In the IID setup, data samples from each class are distributed equally to $M=50$ clients. To simulate non-IID conditions across clients, we follow Hsu et al. [2019]; Nguyen et al. [2023]; Chen et al. [2023] to partition the training set into $M=50$ clients using a Dirichlet distribution with $\alpha\in[0.5,0.1,0.05]$.
Calibration baselines. We compare our method against models trained using Cross-Entropy (Uncal.), representing training without calibration, and others such as Focal Loss (FL) Lin et al. [2017], Label Smoothing (LS) Müller et al. [2019], Brier Score (BS) Brier [1950], MMCE Kumar et al. [2018], FLSD Mukhoti et al. [2020], MbLS Liu et al. [2021], DCA Liang et al. [2020], and MDCA Hebbalaguppe et al. [2022b]. Detailed parameters for each calibration baseline are provided in the Appendix A.1.
FL algorithms and setups. We conduct model calibration on five well-known FL algorithms, including FedAvg McMahan et al. [2017], FedProx Sahu et al. [2018], Scaffold Karimireddy et al. [2019], FedDyn Acar et al. [2021], and FedNova Wang et al. [2020b]. We run each FL algorithm for 100 rounds, evaluating the final global model, with 5 epochs for each local training. We use the SGD optimizer with a learning rate of $10^{-3}$, weight decay of $10^{-4}$, and momentum of $0.9$. For additional details on the training specifics of each algorithm, please see Appendix A.2.
@@
Evaluation metrics. To evaluate performance, we measure the accuracy (%) of the final global model on the test set and assess calibration using standard metrics such as Expected Calibration Error (ECE%) and Static Calibration Error (SCE%). ECE Naeini et al. [2015], a commonly used metric, approximates calibration error (3) by dividing predictions into $I$ bins and calculating a weighted average of the discrepancies between accuracy and confidence across these bins. Predictions are categorized into $I$ interval bins based on their probabilities: $\text{ECE}=\sum_{i=1}^{I}\frac{B_{i}}{N}\left|A_{i}-C_{i}\right|$, where $N$ represents the total number of samples, with weighting based on the proportion of samples within each confidence bin/interval. Each $i$-th bin covers the interval $(\frac{i-1}{I},\frac{i}{I}]$ within the confidence range, with $B_{i}$ indicating the number of samples in the $i$-th bin. $A_{i}=\frac{1}{|B_{i}|}\sum_{j\in B_{i}}\mathbb{I}(\hat{y}_{j}=y_{j})$ calculates the accuracy within bin $B_{i}$, and $C_{i}=\frac{1}{|B_{i}|}\sum_{j:\hat{s}_{j}\in B_{i}}\hat{s}_{j}$ gives the average predicted confidence for samples where $\hat{s}_{j}\in B_{i}$. The recently proposed SCE Nixon et al. [2019]extending ECE for multi-class settings, groups predictions into bins by class probability, computes calibration errors per bin, and averages these across all bins: $\text{SCE}=\frac{1}{K}\sum_{i=1}^{I}\sum_{j=1}^{K}\frac{B_{i,j}}{N}\left|A_{i,j}-C_{i,j}\right|$, where $K$ represents the number of classes, and $B_{i,j}$ denotes the count of $j$-th class samples in the $i$-th bin. $A_{i,j}=\frac{1}{|B_{i,j}|}\sum_{k\in B_{i,j}}\mathbb{I}(j=y_{k})$ measures accuracy, while $C_{i,j}=\frac{1}{|B_{i,j}|}\sum_{k\in B_{i,j}}s_{k}[j]$ gives the average confidence for the $j$-th class in the $i$-th bin. We set the number of bins, $I=20$, for ECE and SCE, and evaluate the final global model.
@@ 5.2 Experimental Results
Can train-time calibration apply to federated learning? We first compare train-time calibration in traditional centralized training versus non-IID federated learning. We train a model on the entire CIFAR-100 training dataset for 100 epochs with various train-time calibration methods, and present the results in Table 6 of Appendix B.1. Results show that most baselines effectively calibrate in centralized settings, maintaining accuracy and exhibiting lower calibration errors. Yet, when these train-time calibration techniques are applied to FL, as depicted in Table 1, these methods often underperform in calibration and accuracy. Notably, methods like DCA and MDCA that includes an auxiliary loss for calibration show robustness in federated settings. They maintain the primary classification objective while penalizing miscalibration, preserving confidence without affecting classification, distinguishing them in FL environments.
@@
Calibration Method FedAvg McMahan et al. [2017] FedProx Sahu et al. [2018] Scaffold Karimireddy et al. [2019] FedDyn Acar et al. [2021] FedNova Wang et al. [2020b]
@@
Uncal. 72.77 9.80 3.21 72.95 4.29 2.20 72.09 5.38 3.81 73.83 11.76 7.29 73.23 6.99 3.63
Focal Lin et al. [2017] 72.61 10.96 3.95 73.62 5.40 3.17 71.63 5.82 4.37 72.64 8.70 5.13 72.48 13.81 7.11
LS Müller et al. [2019] 71.26 13.59 3.62 73.09 6.08 3.49 72.90 4.13 2.67 73.65 7.57 4.94 72.26 5.83 3.53
BS Brier [1950] 71.26 9.98 3.42 72.72 5.81 2.76 71.00 4.59 3.04 72.83 15.06 8.43 71.26 5.77 3.49
MMCE Kumar et al. [2018] 72.64 11.66 3.67 72.64 5.26 2.94 72.46 6.77 4.95 71.95 9.27 6.24 73.99 6.32 5.68
FLSD Mukhoti et al. [2020] 71.10 10.29 4.06 74.23 8.45 4.29 70.39 4.84 3.35 71.95 7.19 4.67 72.35 6.83 6.07
MbLS Liu et al. [2021] 70.16 7.99 3.98 73.42 4.44 2.73 73.11 5.97 4.75 72.58 9.69 5.69 72.38 5.16 3.44
DCA Liang et al. [2020] 72.39 6.95 3.07 73.05 4.07 2.07 72.10 4.59 3.08 73.77 8.63 4.96 73.88 5.73 3.45
NUCFL (DCA+COS) 72.97 5.85 2.99 72.84 4.02 1.89 72.53 3.87 2.58 73.48 7.03 4.35 73.84 4.62 3.04
NUCFL (DCA+L-CKA) 73.73 5.14 2.87 73.16 3.99 1.85 71.95 3.95 2.93 73.49 6.79 4.05 73.87 4.11 2.98
NUCFL (DCA+RBF-CKA) 72.20 6.66 3.10 72.95 4.09 1.95 73.05 3.84 2.41 73.23 6.80 4.03 73.69 4.15 2.94
MDCA Hebbalaguppe et al. [2022b] 71.47 7.69 3.19 72.97 4.11 2.13 72.15 4.99 3.37 73.61 7.51 4.93 72.95 5.08 3.15
NUCFL (MDCA+COS) 72.24 5.24 3.03 72.78 3.82 2.04 72.07 3.96 2.89 74.06 6.99 4.33 73.93 4.28 2.91
NUCFL (MDCA+L-CKA) 73.64 5.20 2.96 73.69 3.79 1.78 72.64 4.01 3.01 73.28 7.10 4.81 72.62 5.15 3.37
NUCFL (MDCA+RBF-CKA) 72.71 6.59 2.98 73.31 3.97 1.90 72.10 4.14 3.14 73.27 6.95 4.15 73.04 4.03 2.79
Table 1: Accuracy (%), calibration measures ECE (%), and SCE (%) of various federated optimization methods with different calibration methods under non-IID ($\alpha=0.5$) scenario on the CIFAR-100 dataset. Underlined values indicate the best calibration across all methods.
Calibration Method FedAvg McMahan et al. [2017] FedProx Sahu et al. [2018] Scaffold Karimireddy et al. [2019] FedDyn Acar et al. [2021] FedNova Wang et al. [2020b]
@@
Uncal. 91.98 4.38 1.83 92.57 4.53 1.74 92.17 5.33 2.44 93.03 5.11 2.39 92.56 4.68 1.76
Focal Lin et al. [2017] 91.71 4.69 1.91 92.06 4.93 2.02 92.25 5.59 2.70 91.13 5.28 2.55 91.42 4.92 2.27
LS Müller et al. [2019] 91.89 3.94 1.55 92.51 4.29 1.63 92.11 5.17 2.31 93.34 4.83 2.26 92.63 4.37 1.70
BS Brier [1950] 91.99 5.06 2.03 92.51 5.53 2.58 89.53 4.03 1.49 90.90 5.18 2.39 90.94 4.23 1.47
MMCE Kumar et al. [2018] 91.65 4.93 2.17 90.66 3.86 1.32 92.25 4.95 1.99 92.80 4.70 1.93 92.94 4.91 2.30
FLSD Mukhoti et al. [2020] 90.97 5.21 2.89 91.73 4.89 1.99 92.91 5.26 2.27 91.90 4.98 2.30 91.45 4.89 2.25
MbLS Liu et al. [2021] 91.93 4.19 1.75 92.43 4.42 1.65 93.05 5.04 2.25 93.10 5.13 2.39 91.74 4.44 1.69
DCA Liang et al. [2020] 92.04 3.99 1.59 92.90 4.23 1.61 92.95 4.38 1.68 93.20 4.52 1.77 92.56 4.20 1.39
NUCFL (DCA+COS) 92.37 3.30 1.35 92.86 3.76 1.37 92.80 4.16 1.49 93.10 4.13 1.55 92.96 4.09 1.17
NUCFL (DCA+L-CKA) 91.89 3.44 1.39 92.56 3.66 1.29 92.23 4.11 1.45 93.37 4.10 1.51 92.56 3.98 1.13
NUCFL (DCA+RBF-CKA) 91.93 3.61 1.43 92.61 3.69 1.33 92.31 4.07 1.44 93.59 4.20 1.59 92.34 3.90 1.13
MDCA Hebbalaguppe et al. [2022b] 91.78 3.66 1.44 93.06 3.98 1.45 93.23 4.45 1.70 93.10 4.58 1.77 92.73 4.17 1.24
NUCFL (MDCA+COS) 92.03 3.75 1.46 92.74 3.71 1.37 92.52 4.09 1.44 92.95 4.25 1.59 92.36 4.05 1.17
NUCFL (MDCA+L-CKA) 92.11 3.19 1.20 91.96 3.84 1.40 92.47 4.17 1.50 92.57 4.16 1.55 91.59 4.01 1.15
NUCFL (MDCA+RBF-CKA) 91.99 3.37 1.35 92.91 3.79 1.47 92.84 4.15 1.50 93.30 4.33 1.62 92.66 3.98 1.14
Table 2: Accuracy (%), calibration measures ECE (%), and SCE (%) of various FL algorithms with different calibration methods under non-IID ($\alpha=0.5$) scenario on the FEMNIST dataset.
@@
Mitigating over/under-confidence. The results in Tables 1 and 2 show that our method improves SCE and ECE over the baselines, but they highlight whether they correct for over-confidence or under-confidence. To clarify this, we present reliability diagrams Degroot & Fienberg [1983]; Niculescu-Mizil & Caruana [2005], which visually compare predicted probabilities to actual outcomes, plotting expected accuracy against confidence. If the model is perfectly calibrated (i.e. if $\mathbb{P}(\hat{y}=y^{*}|s[\hat{y}])=s[\hat{y}]$), then the diagram should plot the identity function. Any deviation from a perfect diagonal represents miscalibration. In Fig. 3, we present reliability histograms for non-IID FL using various calibration methods, where each FL method yields over-confident predictions. This means that, for example, if a model predicts the probability of a class $k$ from $\{1,...,K\}$ as 0.7, we would expect it to be correct in 70 out of 100 cases. However, if the actual number of correct predictions are less than 70, the model is considered over-confident. Such over-confidence can be problematic, especially in critical applications like medical diagnostics, where it could lead to under-treatment and serious health risks. Our methods show a smaller gap (red region) and lower ECE, effectively mitigating over-confidence. For details on how our method also tackles under-confidence, please refer to Appendix B.3.
@@
Effectiveness of our design. We explore model calibration on various FL settings, analyzing the effects of different non-IID degrees $\alpha$, numbers of clients $M$, and percentages of participating clients per round. Fig. 4 shows the accuracy and calibration results of FedAvg across various configurations on the MNIST dataset, where we adjust one variable at a time from the default setting: $M=50$, $\alpha=0.5$, 100% participation. Across all settings, our method consistently demonstrates robust calibration. Importantly, our NUCFL maintains effectiveness in more challenging conditions, such as with lower $\alpha$ values indicating higher non-IID disparity or reduced participant numbers per round. This demonstrates the adaptability of our approach in varied and complex FL environments.
Calibration Method Non-IID FedAvg ($\alpha=0.5$)
@@
NUCFL (DCA+L-CKA) 73.73 5.14 2.87
NUCFL (MDCA+L-CKA) 73.64 5.20 2.96
Reversed NUCFL (DCA+L-CKA) 72.19 10.17 3.88
Reversed NUCFL (MDCA+L-CKA) 71.86 9.64 3.21

### v3

@@ 4.2 Non-Uniform Calibration for Federated Learning (NUCFL)
Motivation: Limitation of uniform penalty. In Equation (5), we promote calibration during federated local training by targeting the confidence-accuracy gap in individual local models via auxiliary calibration loss. The main aim of FL is to train a global model capable of effectively handling the global data distribution. Therefore, we aim for calibration to account for the heterogeneity inherent in FL setups, marking a departure from calibration approaches used in centralized learning. Direct application of auxiliary loss calibration methods without adapting to FL characteristics would result in uniform calibration weights, such that $\beta_{1}=\beta_{2}=...=\beta_{m}$. This leads to local models being calibrated solely to their respective datasets $D_{m}$, with uniformly small weights biasing calibration toward local heterogeneity, and uniformly large weights potentially neglecting accuracy improvements from classification. Such discrepancies can result in a global model that, while comprising well-calibrated local models, does not achieve optimal performance across the broader distribution. This fixed calibration penalty approach, therefore, risks neglecting the broader global calibration needs necessary for optimal performance across the entire global distribution, e.g., as in Fig. 1(d).
@@
Non-uniform penalty design (NUCFL). Given that each client receives the global model from the previous round as their starting point for each new federated round, this offers an opportunity to indirectly adjust local calibration to mirror global model characteristics. To leverage this, we propose Non-Uniform Calibration for Federated Learning (NUCFL), which provides a dynamic calibration penalty to local models to better reflect global calibration needs. Drawing inspiration from the concept of similarity characterization in FL techniques (Sahu et al., 2018; Tan et al., 2023), we hypothesize that a local model closely resembling the global model is likely to represent global characteristics well, suggesting that the penalty appropriately reflects the calibration needs of the global model. Conversely, a dissimilar/heterogeneous local model suggests a focus on local objectives (e.g., to improve accuracy) at the expense of global alignment. By adjusting penalties at each local epoch based on the similarity between local and global models, NUCFL ensures that local training enhances both the calibration and accuracy of the global model. As illustrated in Fig. 2, we specifically set
@@
where $\delta^{(t-1)}$ is the accumulated gradient from the previous round $t-1$, representing the received global model, while $\delta_{m}^{(t,e)}={w}^{(t-1)}-{w}^{(t,e)}_{m}$ calculates the gradient for client $m$ at local epoch $e$ during round $t$. The function sim($\cdot$) measures similarity between local and global models, generating values between 0 and 1. Specifically, it can use general cosine similarity or more advanced measures like Centered Kernel Alignment (CKA) (Kornblith et al., 2019).
For instance, if the local model is similar to the global model, indicating alignment in their characteristics, a higher calibration penalty is applied to the local model to address global calibration needs. Otherwise, if the similarity score is low, a smaller penalty is applied, focusing on accuracy improvement instead. Our non-uniform penalty method tailored to FL, strategically improving model calibration and accuracy. Additionally, our NUCFL requires no extra hyperparameter tuning, as $\beta_{m}$ is automatically determined by (8) based on the given similarity function.
Remark 1. NUCFL can integrate any auxiliary-based calibration loss and adapt to any FL algorithm, independent of aggregation method or local loss function. To our knowledge, this is the early work to specifically address model calibration across various FL algorithms. Similar to most calibration research (Lin et al., 2017; Müller et al., 2019; Brier, 1950; Kumar et al., 2018; Mukhoti et al., 2020; Liu et al., 2021; Liang et al., 2020; Hebbalaguppe et al., 2022b; a), our work is an empirical study that explores the practical impacts of our methods. This research area remains challenging, particularly in providing theoretical guarantees for calibration methods, even in centralized settings; we instead substantiate the effectiveness of our method through extensive experiments. We will demonstrate how our strategic approach effectively calibrates FL in Section 5.2.
@@ 5.1 Experimental Setup
Dataset and model. We conduct experiments using four image classification datasets commonly utilized in FL research (Caldas et al., 2018; McMahan et al., 2017; Mohri et al., 2019): MNIST (LeCun et al., 1998), FEMNIST (Cohen et al., 2017), CIFAR-10 (Krizhevsky, 2009), and CIFAR-100 (Krizhevsky, 2009). For the MNIST and FEMNIST datasets, we use a CNN (LeCun & Bengio, 1998) with two 5x5 convolution layers, each followed by 2x2 max pooling, and fully connected layers with ReLU activation. We use AlexNet (Krizhevsky et al., 2012) with five convolutional layers and ResNet-34 (He et al., 2015) for the CIFAR-10 and CIFAR-100 datasets, respectively.
FL data distribution. In the IID setup, data samples from each class are distributed equally to $M=50$ clients. To simulate non-IID conditions across clients, we follow (Hsu et al., 2019; Nguyen et al., 2023; Chen et al., 2023) to partition the training set into $M=50$ clients using a Dirichlet distribution with $\alpha=0.5$. Results with different $\alpha$ are reported in the Appendix B.2.
Calibration baselines. We compare our method against models trained using Cross-Entropy (Uncal.), representing training without calibration. We adapt centralized train-time calibration methods, such as Focal Loss (FL) (Lin et al., 2017), Label Smoothing (LS) (Müller et al., 2019), Brier Score (BS) (Brier, 1950), MMCE (Kumar et al., 2018), FLSD (Mukhoti et al., 2020), MbLS (Liu et al., 2021), DCA (Liang et al., 2020), and MDCA (Hebbalaguppe et al., 2022b), to the FL setting based on (5) and (6) to evaluate their performance. We also compare our method with FedCal (Peng et al., 2024), a state-of-the-art method designed for post-hoc calibration in FL. Detailed parameters for each calibration baseline are provided in Appendix A.1.
FL algorithms and setups. We conduct model calibration on five well-known FL algorithms, including FedAvg (McMahan et al., 2017), FedProx (Sahu et al., 2018), Scaffold (Karimireddy et al., 2019), FedDyn (Acar et al., 2021), and FedNova (Wang et al., 2020b). We run each FL algorithm for 100 rounds, evaluating the final global model, with 5 epochs for each local training. We use the SGD optimizer with a learning rate of $10^{-3}$, weight decay of $10^{-4}$, and momentum of $0.9$. For additional details on the training specifics of each algorithm, please see Appendix A.2.
@@
Evaluation metrics. To evaluate performance, we measure the accuracy (%) of the final global model on the test set and assess calibration using commonly used metrics such as Expected Calibration Error (ECE%) and Static Calibration Error (SCE%). ECE (Naeini et al., 2015) approximates calibration error (3) by dividing predictions into $I$ bins and calculating a weighted average of accuracy-confidence discrepancies: $\text{ECE}=\sum_{i=1}^{I}\frac{B_{i}}{N}\left|A_{i}-C_{i}\right|$, where $N$ represents the total number of samples, with weighting based on the proportion of samples within each confidence bin/interval. Each $i$-th bin covers the interval $(\frac{i-1}{I},\frac{i}{I}]$ within the confidence range, with $B_{i}$ indicating the number of samples in the $i$-th bin. $A_{i}=\frac{1}{|B_{i}|}\sum_{j\in B_{i}}\mathbb{I}(\hat{y}_{j}=y_{j})$ calculates the accuracy within bin $B_{i}$, and $C_{i}=\frac{1}{|B_{i}|}\sum_{j:\hat{s}_{j}\in B_{i}}\hat{s}_{j}$ gives the average predicted confidence for samples where $\hat{s}_{j}\in B_{i}$. The recently proposed SCE (Nixon et al., 2019), extending ECE for multi-class settings, groups predictions into bins by class probability, computes calibration errors per bin, and averages these across all bins: $\text{SCE}=\frac{1}{K}\sum_{i=1}^{I}\sum_{j=1}^{K}\frac{B_{i,j}}{N}\left|A_{i,j}-C_{i,j}\right|$, where $K$ represents the number of classes, and $B_{i,j}$ denotes the count of $j$-th class samples in the $i$-th bin. $A_{i,j}=\frac{1}{|B_{i,j}|}\sum_{k\in B_{i,j}}\mathbb{I}(j=y_{k})$ measures accuracy, while $C_{i,j}=\frac{1}{|B_{i,j}|}\sum_{k\in B_{i,j}}s_{k}[j]$ gives the average confidence for the $j$-th class in the $i$-th bin. We set the number of bins, $I=20$, for ECE and SCE, and evaluate the final global model. We performed all the experiments on four random seeds and reported the average performance along with the standard deviation.
@@ 5.2 Experimental Results
Can train-time calibration apply to federated learning? We first compare train-time calibration in traditional centralized training and non-IID FL. Using the entire CIFAR-100 dataset, we train models for 100 epochs with various calibration methods, presenting the results in Table 6 of Appendix B.1. Most methods effectively maintain accuracy and reduce calibration errors in centralized settings. Yet, as shown in Table 1, these methods often underperform in calibration and accuracy when applying to FL. Notably, methods like DCA and MDCA, which include an auxiliary loss, show robustness in federated settings. They balance classification accuracy and calibration, preserving confidence while penalizing miscalibration, making them stand out in FL environments.
@@
Calibration Method FedAvg FedProx Scaffold FedDyn FedNova
@@
Uncal. 61.34 10.52 3.61 61.88 9.37 3.58 62.08 11.42 3.82 62.39 12.53 3.84 63.01 11.45 3.82
Focal (Lin et al., 2017) 60.59 12.88 3.74 62.07 11.45 3.79 61.39 13.28 4.88 60.21 11.22 3.81 61.15 13.21 4.71
LS (Müller et al., 2019) 59.35 15.61 6.28 62.23 14.37 6.14 62.15 11.69 4.05 61.34 15.19 6.58 63.05 16.03 6.89
BS (Brier, 1950) 61.20 11.32 3.80 60.43 10.15 3.61 60.39 10.92 3.80 62.17 13.81 4.91 60.44 11.39 4.06
MMCE (Kumar et al., 2018) 60.00 13.41 4.93 60.11 11.32 3.77 61.03 11.37 3.83 61.35 12.93 3.95 61.28 12.55 3.90
FLSD (Mukhoti et al., 2020) 58.71 11.51 3.92 59.28 10.66 3.81 60.49 13.61 5.15 60.94 11.83 4.04 60.49 13.81 4.88
MbLS (Liu et al., 2021) 59.62 9.42 3.55 60.39 11.37 3.84 61.38 11.99 4.07 62.30 14.95 6.50 62.93 13.66 4.69
DCA (Liang et al., 2020) 61.24 7.69 3.24 61.93 8.74 3.40 62.11 9.25 3.55 62.93 10.17 3.75 63.15 9.27 3.55
NUCFL (DCA+COS) 61.88 6.21 3.11 62.38 8.15 3.35 62.17 8.88 3.50 62.81 9.29 3.54 63.24 8.85 3.51
NUCFL (DCA+L-CKA) 62.05 6.14 3.07 62.31 8.04 3.30 62.25 8.41 3.45 62.94 9.14 3.52 63.27 8.52 3.43
NUCFL (DCA+RBF-CKA) 61.59 6.19 3.11 61.89 8.17 3.35 61.94 8.52 3.45 62.84 9.21 3.55 63.17 8.01 3.30
MDCA (Hebbalaguppe et al., 2022b) 61.03 7.71 3.29 62.00 8.21 3.37 62.23 9.04 3.51 62.84 10.24 3.72 63.29 10.00 3.71
NUCFL (MDCA+COS) 62.00 6.38 3.14 61.93 7.94 3.29 62.17 8.31 3.40 62.91 9.33 3.56 63.14 9.16 3.58
NUCFL (MDCA+L-CKA) 62.17 6.25 3.11 62.03 7.88 3.25 62.22 8.30 3.40 62.88 9.19 3.53 63.14 9.03 3.51
NUCFL (MDCA+RBF-CKA) 61.54 6.20 3.09 61.79 8.02 3.29 62.15 8.42 3.45 62.65 9.24 3.55 63.22 8.59 3.47
Table 1: Accuracy (%), calibration measures ECE (%), and SCE (%) of various federated optimization methods with different calibration methods under non-IID scenario on the CIFAR-100 dataset. Values in boldface represent the best calibration provided by our method for the auxiliary calibration method, and underlined values indicate the best calibration across all methods. Averaged performances are reported; complete results with standard deviation are included in the Appendix.
Calibration Method FedAvg FedProx Scaffold FedDyn FedNova
@@
Uncal. 90.95 4.17 1.66 91.45 4.61 1.95 91.28 5.02 2.51 92.75 5.24 2.60 92.22 4.92 2.44
Focal (Lin et al., 2017) 90.63 4.77 1.96 90.00 5.39 2.77 91.13 5.38 2.72 91.15 5.84 2.91 91.13 5.39 2.63
LS (Müller et al., 2019) 91.07 3.75 1.62 90.37 4.52 1.90 91.33 5.18 2.64 93.08 5.19 2.63 92.04 4.95 2.46
BS (Brier, 1950) 91.48 5.19 2.65 90.98 5.42 2.80 88.72 4.19 1.81 91.39 5.37 2.62 90.38 4.65 1.92
MMCE (Kumar et al., 2018) 90.22 4.85 2.30 90.12 4.01 1.79 91.44 5.07 2.55 92.11 4.93 2.49 91.35 5.08 2.49
FLSD (Mukhoti et al., 2020) 90.02 4.93 2.95 90.39 4.99 2.04 92.88 5.19 2.58 91.17 5.61 2.84 91.22 5.23 2.63
MbLS (Liu et al., 2021) 90.62 4.27 1.99 91.49 4.79 1.94 92.87 5.39 2.60 92.06 5.44 2.75 90.39 4.61 1.92
DCA (Liang et al., 2020) 91.84 3.61 1.52 92.03 4.25 1.61 92.04 4.44 1.82 93.08 4.61 1.92 92.37 4.31 1.71
NUCFL (DCA+COS) 91.77 3.52 1.49 91.95 3.61 1.35 92.42 4.39 1.77 93.22 4.20 1.65 92.25 4.13 1.68
NUCFL (DCA+L-CKA) 91.63 3.52 1.47 92.10 3.60 1.30 92.19 4.09 1.60 93.45 4.22 1.65 92.33 4.02 1.54
NUCFL (DCA+RBF-CKA) 91.74 3.49 1.40 91.99 3.77 1.35 91.74 4.15 1.62 92.95 4.34 1.75 92.41 4.11 1.68
MDCA (Hebbalaguppe et al., 2022b) 91.64 3.75 1.53 92.17 4.42 2.05 92.95 4.61 1.90 93.19 4.71 1.93 92.05 4.62 1.82
NUCFL (MDCA+COS) 91.29 3.61 1.44 91.95 3.95 1.77 93.07 4.14 1.62 92.88 4.41 1.80 92.45 4.32 1.70
NUCFL (MDCA+L-CKA) 91.97 3.28 1.28 92.20 3.88 1.51 92.77 4.20 1.80 93.11 4.31 1.75 91.99 4.28 1.67
NUCFL (MDCA+RBF-CKA) 91.35 3.56 1.40 92.21 4.00 1.77 93.04 4.19 1.79 93.04 4.40 1.80 92.39 4.17 1.67
Table 2: Average performance of each algorithm under non-IID scenario on the FEMNIST dataset. Complete results with standard deviation are included in the Appendix.
@@
Mitigating over/under-confidence. The results in Tables 1 and 2 show that our method improves SCE and ECE over the baselines, but they highlight whether they correct for over-confidence or under-confidence. To clarify this, we present reliability diagrams (Degroot & Fienberg, 1983; Niculescu-Mizil & Caruana, 2005), which visually compare predicted probabilities to actual outcomes, plotting expected accuracy against confidence. If the model is perfectly calibrated (i.e. if $\mathbb{P}(\hat{y}=y^{*}|s[\hat{y}])=s[\hat{y}]$), then the diagram should plot the identity function. Any deviation from a perfect diagonal represents miscalibration. In Fig. 3 and 4, we present reliability histograms for non-IID FL using various calibration methods, where each FL method yields over-confident predictions. For example, if a model predicts the probability of a class $k$ from $\{1,...,K\}$ as 0.7, it is expected to be correct 70 out of 100 cases; if the actual number of correct predictions is less than 70, the model is considered over-confident. Such over-confidence can be problematic in critical applications like medical diagnostics, potentially leading to under-treatment and serious health risks. Our method shows a smaller gap (red region) and lower ECE, effectively mitigating over-confidence. For details on how our method also tackles under-confidence, please refer to Appendix B.3.
@@
Effectiveness of our design. We explore model calibration on various FL settings, analyzing the effects of different non-IID degrees $\alpha$, numbers of clients $M$, and percentages of participating clients per round. Fig. 5 shows the accuracy and calibration results of FedAvg across various configurations on the CIFAR-100 dataset, where we adjust one variable at a time from the default setting: $M=50$, $\alpha=0.5$, 100% participation. Across all settings, our method consistently shows robust calibration. Notably, NUCFL remains effective under challenging conditions, such as with lower $\alpha$ values (higher non-IID disparity) or fewer participant per round. Additional results under diverse conditions, including an even smaller number of partial participants under various data distributions, are provided in Appendix B.4, highlighting the adaptability of our approach in complex FL environments.
Figure 3: Reliability diagrams for non-IID FedAvg with different calibration methods using the CIFAR-100 dataset. The lower ECE and smaller gap (red region) show the effectiveness of our method.
Figure 4: Reliability diagrams for non-IID FedAvg using the CIFAR-10 dataset.
Figure 5: Comparison of confidence calibration across different FL settings using the CIFAR-100 dataset.
Calibration Method Non-IID FedAvg
@@
NUCFL (DCA+L-CKA) 62.05 6.14 3.07
NUCFL (MDCA+L-CKA) 62.17 6.25 3.11
Reversed NUCFL (DCA+L-CKA) 60.36 13.89 4.98
Reversed NUCFL (MDCA+L-CKA) 61.77 16.11 6.87